OpenStudy (anonymous):

The quadratic equation q(x^2+2x+1)=x(2x+3) has two different roots.Determine the range of values of q.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (only4u):

for different the discriminant should be greater than zero .. solving it u will get ur answer

OpenStudy (anonymous):

\[qx^2-2x^2-2qx-3x+q=0\]

OpenStudy (only4u):

(q-2)x^2 +2qx +(q-3) =0

OpenStudy (only4u):

find the discriminant

OpenStudy (anonymous):

\[(q-2)x^2-x(2q+3)+q=0\]

OpenStudy (anonymous):



OpenStudy (only4u):

nice .. and sorry for the mistake

OpenStudy (only4u):

yes .. u r going right

OpenStudy (anonymous):